Studies have shown that 3 out of 10 thirteen year olds have a key to their
home. Of those who have a key, 22% are alone at home after school.
a) What is the probability that a 13 year
old does not have a key to their home?
b) What percent of 13 year olds have a key and someone else at home after
school?

Thanks

Answer by ikleyn(52754)

.

(a)  P = 1 - 0.3 = 0.7.                                  ANSWER



(b)  P = 0.3*(1 - 0.22) = 0.3*0.78 = 0.234 = 23.4%.      ANSWER

Quite obvious . . .